A 5.5-kg rifle fires an 11 g bullet with a velocity of 1000 m/s. Question: if the shooter holds the rifle firmly against his shoulder,what will be the recoil velocity? Assume the mass of the shooter is 120 kg.
1
Expert's answer
2017-11-28T15:26:06-0500
Initial momentum:
P_i=mv=(0.011)(1000) kgm/s
Initial momentum:
P_f=(M+m^' )V.
From the conservation of momentum:
(M+m^' )V=mv
The recoil velocity is
V=m/((M+m^' ) ) v=(0.011)(1000)/(120+5.5)=0.09 m/s.
Answer: 0.09 m/s.